OpenStudy (anonymous):

one has been supplied with a 20mL of a 3.5M NaOH solution. If one requires a 0.35M NaOH solution, what volume of diluent would be need to be added to the provided sample

OpenStudy (anonymous):

oh Chemistry,, just use M1/V1=M2/V2 when you found the second volume, subtract it with the first volume to get the amount volume needed to be added

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

my over all answer has to be 180mL and i dont get how to even get 180mL

OpenStudy (anonymous):

Hey @hena5 \[\frac{ M _{1} }{ V _{1} }=\frac{ M _{1} }{ V _{1} }\] M1 = 3.5 M M2 = 0.35 M V1 = 20 mL V2 = ? just take it to the equation and you will eventually get 200 mL for V2 so first we already have 20 mL, then to get it to be 200 mL, we need to add 180 mL Tada!

OpenStudy (anonymous):

thanks @edwinls9

OpenStudy (anonymous):

given a stock solution that is 25.00g/L CuSO4.5H2O what is the concentration of this stock in milligrams per mililiters?

OpenStudy (anonymous):

1 gram is 1000 miligram 1 liter is 1000 mililiter 25.00 x1000 miligram / 1000 mililiter = 25.00 miligram/mililiter still the same